A Complete Bibliography of Publications in Journal of Computational and Applied Mathematics (2025–2029)

Nelson H. F. Beebe University of Utah

Department of Mathematics, 110 LCB

155 S 1400 E RM 233 Salt Lake City, UT 84112-0090

USA

Tel: +1 801 581 5254 FAX: +1 801 581 4148

E-mail: beebe@math.utah.edu, beebe@acm.org,
beebe@computer.org (Internet)
WWW URL: https://www.math.utah.edu/~beebe/

21 September 2024 Version 1.03

Title word cross-reference

k [DJ25, JFP25]. l^{2} [Xie25]. l_{1/2} [CLJX25]. R [CdlHG25]. n [Ran25]. n: G [DJ25]. W̃_{2}(m, m − 1) [SHK25]. w [AAN25].

-dimensional [Ran25]. -norm [Xie25]. -out-of- [DJ25]. -record [JFP25].

1 [Ano25a, Ano25b]. 15 [Ano25c].

2025 [Ano25h, Ano25a, Ano25c, Ano25b].

accuracy [BAPPF25]. accurate [YWW25]. ADAM [SWWY25]. ADAM-type [SWWY25]. adaptive [ZMG25]. adaptivity [KS25b]. additive [Ran25]. adversarial [AGGM25]. affinely [Sah25]. algorithm [CLJX25, FA25, GLTW25, HBG25, JJM25, SQLZ25, SWWY25]. Allen [HKH^{+}25]. Almost [CCLH25]. American [ZZ25]. Analysis [DLR25, JFP25, ADY25, LGDN25, SX25, SGRS25, TC25, ZLS25, ZC25, ZFCM25]. anisotropic [WX25]. annular [YWW25]. Application [FKVZ25, BRPP25]. approach [HBG25, SGRS25]. Approaches [ZFCM25]. Approximate [KSC25, ZNL25]. approximates [Xie25]. approximating [SHK25]. Approximation [KCM25, CA25, CdlHG25, Öza25, TC25, YC25]. approximations [GHZ25]. arbitrary [LC25, PY25]. argument [FA25]. argument-based [FA25]. arrangement [MT25]. arrival [BC25]. Asian [AAN25]. assets [ZFCM25]. assimilation [DLR25]. asymptotic [ZC25]. augmented [ZNL25]. authentic [LGDN25].

B [LWL^{+}25]. B-spline [LWL^{+}25]. backward [CP25]. bank [ZFCM25]. based [FA25, GLTW25, JFP25, JJM25, LWL^{+}25, MS25, WCS25, You25]. basis [KCM25, WYZ25]. Bayesian [YXM25]. Bessel [LC25, ZC25]. between [Xie25]. BINO [YXM25]. bioconvective [UX25]. Block [GH25, BKST25, CLJX25, KSC25]. block-by-block [KSC25]. blurred [JJM25]. Board [Ano25d, Ano25e, Ano25f, Ano25g]. book [DG25]. bounded [FKVZ25]. Bounds [CP25, BRPP25]. Brinkman [JLC25]. Brownian [AAN25].

Cahn [HKH^{+}25]. calculus [Öza25]. Can [NT25]. chain [HBG25]. chaotic [ADY25]. class [WX25, WCS25]. cold [MHWY25]. collected [DJ25]. collocation [GHD25, HRST25]. complementarity [WCS25]. complex [FKVZ25, FA25]. components [DJ25]. composite [LGDN25]. compressed [CLJX25]. Computational [UX25]. computations [ZWSC25]. Computing [JVW25]. conditions [Ran25, Xie25]. conduction [OTD^{+}25]. conductivity [Sai25]. conformable [GLTW25]. connectivity [ZLY25]. conserving [ZMG25]. constrained [CLJX25, JJM25]. constraint [GHZ25]. constraints [ZLY25, ZFCM25]. Construction [YC25]. continuity [JJM25]. continuous [DLR25]. continuously [SQLZ25]. contractive [KIF^{+}25]. control [GHZ25, GLTW25, SKM25, SWWY25, ZFCM25]. controllability [SX25]. controlled [KIF^{+}25]. convergence [LGDN25]. converging [KS25a]. Convex [AGGM25]. copositive [HLS25]. core [BPT25]. COS [ACO25]. cost [AAN25]. coupled [ZMG25]. credit [HBG25]. crystal [LLZH25, WX25]. Cubic [LWL^{+}25]. curves [FKVZ25]. cyclic [BRPP25].

Darcy [JLC25, You25]. Darcy-transport [JLC25]. data [DJ25, DLR25, SQLZ25]. decomposition [CWX25, LC25, SGRS25]. deep [SSM25, ZLS25]. DEIM [CWX25]. delay [GHD25, HRST25]. dendritic [WX25]. dependent [OTD^{+}25]. derivatives [GLTW25]. Difference [AGGM25, HKH^{+}25]. differentiable [SQLZ25]. differential [AFA25, KSC25, LLG25, Ran25, SX25, SKM25]. diffusion [ACO25, PY25]. dimensional [Ran25, ZMG25]. discontinuous [JLC25]. discrete [ACO25, CCLH25, GH25, MHWY25]. discrete-time [CCLH25, MHWY25]. discretization [SWWY25, SSM25, ZLS25]. discretization-invariant [ZLS25]. distance [Xie25]. distributed [HRST25]. distributed-order [HRST25]. distribution [CP25, JFP25]. distributions [ACO25]. Domain [KS25a]. domains [YWW25, ZMG25]. driven [Ran25]. dual [CQ25]. dynamics [ADY25].

Editorial

[Ano25d, Ano25e, Ano25f, Ano25g].

Efficient

[CM25, ZNL25, CLJX25, GHZ25, GHD25, SGRS25, SWWY25, WX25, YWW25]. eigenvalue [HLS25, Sah25, TC25]. eigenvalues [Sah25]. elastic [LWL^{+}25, LGDN25]. element [GHZ25, OTD^{+}25]. elements [NT25, SWWY25]. elliptic [GHZ25, SWWY25, YWW25]. Emden [SGRS25]. empirical [GH25]. Energy [ZMG25, LLZH25]. Energy-conserving [ZMG25]. energy-stable [LLZH25]. Enhancing [CA25]. environment [AAN25]. epidemic [KS25a]. equation [AFA25, BKST25, HKH^{+}25, KIF^{+}25, LLZH25, SSM25, ZZ25]. equations [BPT25, GHD25, HRST25, JJM25, KSC25, LLG25, MS25, PY25, Ran25, SX25, SQLZ25, SKM25, SGRS25, YWW25]. equilibrium [PY25]. equilibrium-diffusion [PY25]. error [CA25, TC25, Xie25]. estimates [CA25]. Estimating [BC25]. estimation [LLG25]. Euler [Ran25, SKM25]. Eulerian [PY25]. evaluation [LC25]. even [DLR25]. example [OTD^{+}25, ZFCM25]. exciting [ACO25]. explicit [KS25a, Ran25]. exponent [MS25]. exponential [CCLH25, CM25]. extension [ACO25, CQ25, KS25b, ZLS25]. extrapolation [LC25].

facility [MHWY25]. factor [LLG25]. factorization [ZNL25]. failure [DJ25]. February [Ano25h]. feedforward [YC25]. Feller [CdlHG25]. field [LLZH25, MT25, WX25]. Fields [BAPPF25]. filled [SQLZ25]. finance [HBG25]. finding [FA25]. finite [GHZ25, HKH^{+}25, NT25, PY25, SWWY25]. fitting [SQLZ25]. fixed [ADY25]. flow [UX25, You25]. fluid [ADY25]. formula [CA25]. formulas [SHK25]. formulation [ZZ25]. Forward [YXM25, CP25]. four [CM25]. fourth [HKH^{+}25, SGRS25]. fourth-order [HKH^{+}25, SGRS25]. Fowler [SGRS25]. FPMPE [ZWSC25]. fractal [KCM25]. fractional [ADY25, ABH^{+}25, AAN25, GLTW25, HRST25, KSC25, Öza25, SKM25]. fractional-order [GLTW25]. fractured [ABH^{+}25]. framework [KS25b]. free [SQLZ25]. function [MPB25, Öza25, SQLZ25, ZC25]. functions [KCM25, LC25, OTD^{+}25, SHK25, YC25]. fundamental [KC25]. fuzzy [AAN25].

Galerkin [JLC25, ZMG25]. Gaussian [BAPPF25, Ran25]. generalized [Ran25, ZZ25]. genuine [CQ25].

Geometric [AAN25, DJ25]. global [FA25]. GMsFEM [SSM25]. Gompertz [JFP25]. Gordon [ZMG25]. grade [UX25]. gradient [GLTW25]. gradient-based [GLTW25]. graph [SNZ25]. graphs [BAPPF25]. gravitational [Sai25]. Gregory [ZBK25]. growth [WX25].

Hawkes [ACO25]. heat [ADY25, OTD^{+}25]. Hermite [ZMG25]. Hessenberg [BKST25]. Heston [ACO25]. heterogeneous [ABH^{+}25, DJ25, UX25]. HH [BRPP25]. high [BAPPF25, CWX25, WX25]. high-order [WX25]. Higher [LGDN25]. Higher-order [LGDN25]. Highly [SGRS25, ZC25]. Hilbert [SHK25]. hodograph [FKVZ25]. homogeneous [UX25]. HOMS [LGDN25]. HOMS-PINN [LGDN25]. horizontal [WCS25]. hydrodynamics [PY25]. hyperplane [WYZ25].

identity [BRPP25]. IDFPM [JJM25]. IDFPM-based [JJM25]. image [JJM25]. imaging [MT25]. impact [DG25]. implementation [BPT25]. implicit [WCS25]. Improved [WCS25, HBG25]. impulses [CCLH25]. inclusion [HLS25]. index [JFP25]. induced [CWX25]. inequalities [BRPP25]. inference [DJ25]. infinite [DLR25, LC25]. Influence [Sai25]. information [BRPP25]. informed [LGDN25]. inhomogeneous [TC25]. initial [Ran25]. instabilities [FA25]. Integral [MPB25, GHZ25, GHD25, KIF^{+}25, MS25, ZZ25]. integrals [LC25, SHK25, ZC25]. integration [KS25b, MS25]. integro [AFA25, SKM25]. integro-differential [AFA25, SKM25]. interpolation [CA25, GH25, OTD^{+}25]. interpolatory [CWX25]. invariant [ZLS25]. Inverse [Sah25, CQ25, YXM25]. Inversion [YXM25]. investigate [FA25]. involving [LC25]. irregular [YWW25]. isotropic [TC25]. iteration [WCS25, ZWSC25]. iterative [GHZ25, ZLY25].

January [Ano25a, Ano25c]. Jensen [BRPP25]. jump [ACO25]. jump-diffusion [ACO25].

Kaczmarz [WYZ25]. KDV [BPT25]. kernel [Öza25]. kind [GHD25]. kinetic [FA25]. Klein [ZMG25]. Kutta [KS25b].

L [CWX25]. L-DEIM [CWX25]. Lagrangian [PY25, ZNL25]. Lane [SGRS25]. large [DLR25]. layers [ADY25]. learn [NT25]. learning [SSM25]. Leffler [Öza25]. length [BC25]. level [KC25]. liabilities [ZFCM25]. lifetime [JFP25]. lifetimes [DJ25]. LightGBM [HBG25]. Lightning [SNZ25]. like [CA25]. likelihood [DJ25, LLG25]. limit [PY25]. linear [CLJX25, GHZ25, LLZH25, SX25, ZFCM25]. Lipschitz [JJM25]. local [Xie25, You25]. low [BKST25]. low-rank [BKST25]. Lucas [LC25].

management [ZFCM25]. mapping [YWW25]. mappings [KIF^{+}25]. March [Ano25b]. Markov [BAPPF25]. matched [OTD^{+}25]. matching [SNZ25]. material [UX25]. materials [LGDN25]. mathematical [SGRS25]. matrices [CQ25, Sah25]. matrix [WCS25, ZNL25]. Maximum [DJ25, LLG25]. measure [SX25]. mechanism [MHWY25]. media [ABH^{+}25, You25]. medium [Sai25, TC25]. mesh [You25]. method [AFA25, ACO25, BKST25, GHD25, HKH^{+}25, JLC25, KC25, KSC25, LWL^{+}25, LGDN25, MS25, OTD^{+}25, SKM25, WYZ25, YWW25, ZLY25, ZMG25, ZWSC25, ZNL25]. methods [CM25, GH25, KS25b, LC25, WX25, WCS25]. metric [FKVZ25, KIF^{+}25]. MFPIO [ZWSC25]. minimizers [Xie25]. Mittag [Öza25]. Mittag-Leffler [Öza25]. mKDV [BPT25]. model [ABH^{+}25, ACO25, KS25a, WX25, ZZ25]. modeling [MHWY25, YXM25]. modification [FKVZ25]. modified [BKST25]. modulus [WCS25]. modulus-based [WCS25]. moments [CP25]. Montgomery [BRPP25]. Moore [CQ25]. motion [AAN25]. Multi [KC25, BPT25, KSC25, LGDN25, LLG25, ZMG25]. multi-core [BPT25]. multi-dimensional [ZMG25]. multi-factor [LLG25]. Multi-level [KC25]. multi-scale [LGDN25]. multi-term [KSC25]. multilevel [You25]. multilinear [ZWSC25]. Multiscale [ABH^{+}25]. multivariate [Öza25, Xie25].

near [MT25]. near-field [MT25]. network [LGDN25, YC25]. networks [NT25, ZLS25]. Neural [YXM25, LGDN25, NT25, YC25]. neutral [AFA25]. noises [Ran25]. Non [DG25, SX25]. non-linear [SX25]. Non-uniformly [DG25]. nonconvex [CLJX25, Xie25]. nonlinear [HRST25, JJM25, SQLZ25, ZMG25]. nonlocal [Sai25]. nonnegative [ZNL25]. norm [Xie25]. Novel [GHZ25, CA25]. nudging [DLR25]. Numerical [BPT25, CdlHG25, ZBK25, AFA25, JVW25, ZFCM25]. Nuttall [MPB25]. Nyström [MS25].

observations [BC25]. online [SSM25]. Operator [YXM25, ZLS25]. operators [CdlHG25, Öza25, YC25]. Optimal [MT25, SHK25, GHZ25, GLTW25, SKM25, SGRS25, SWWY25, ZFCM25]. optimization [GLTW25, MHWY25]. optimized [HBG25]. option [AAN25]. options [ACO25, ZZ25]. order [AFA25, CM25, CWX25, DG25, GLTW25, HKH^{+}25, HRST25, LLZH25, LGDN25, LC25, Ran25, SGRS25, WX25]. order-book [DG25]. oscillating [SHK25]. oscillatory [ZC25].

PageRank [ZWSC25]. parallelization [BPT25]. parameter [SQLZ25]. parameter-free [SQLZ25]. parameters [DJ25, DLR25]. Pareto [HLS25]. paths [BAPPF25]. Penrose [CQ25]. performance [JFP25]. periodic [SHK25]. phase [LLZH25, WX25, ZFCM25]. phase-field [WX25]. physics [LGDN25]. physics-informed [LGDN25]. PINN [LGDN25]. planar [FKVZ25]. plasma [FA25]. point [ADY25]. poles [FA25]. polyharmonic [KC25]. polynomial [Sah25]. polynomials [FKVZ25]. porous [You25]. positivity [PY25]. positivity-preserving [PY25]. posteriori [You25]. posteriori-based [You25]. power [AAN25]. PR [CLJX25]. Prediction [SSM25, HBG25]. prescribed [Sah25]. preserving [KS25a, PY25]. price [DG25]. pricing [AAN25, ACO25, ZZ25]. problem [ABH^{+}25, JLC25, OTD^{+}25, Sah25, TC25, WYZ25, You25]. problems [CLJX25, GHZ25, GLTW25, JJM25, KC25, LGDN25, SQLZ25, SKM25, SWWY25, WCS25, YXM25]. process [ACO25, CP25]. product [LC25, MS25]. programming [AGGM25]. programs [CLJX25]. Projective [KS25b]. propagation [LWL^{+}25]. pursuit [WYZ25]. Pythagorean [FKVZ25]. Pythagorean-hodograph [FKVZ25].

quadratic [JVW25, Xie25]. quadrature [SHK25, ZBK25, ZC25]. Queue [ACO25, BC25].

radial [KCM25]. radiation [PY25]. Random [BAPPF25, SWWY25]. range [JVW25]. rank [BKST25]. rate [YC25]. rates [BC25]. reactions [UX25]. record [JFP25]. rectangular [KIF^{+}25]. recurrence [CP25]. reduction [ABH^{+}25, Xie25]. refinement [You25]. regime [ZZ25]. regime-switching [ZZ25]. region [Xie25]. regularized [CLJX25, WYZ25]. relative [BAPPF25]. renewal [CP25]. repair [MHWY25]. repairable [MHWY25]. representations [MPB25]. restoration [JJM25]. retrial [MHWY25]. Richards [SSM25]. Riesz [CdlHG25]. risk [AAN25, HBG25]. Rosenbrock [CM25]. rotating [Sai25]. Runge [KS25b].

sampled [DG25]. SAV [ZMG25]. scale [LGDN25]. scheme [HRST25, KS25a, LLZH25, PY25, Ran25, ZMG25, ZNL25]. schemes [GHZ25]. search [HBG25]. Second [AFA25, LLZH25, Ran25]. Second-order [AFA25, LLZH25, Ran25]. sections [OTD^{+}25]. self [ACO25]. self-exciting [ACO25]. sensing [CLJX25]. sensor [MT25]. sequence [BPT25]. series [MPB25]. service [BC25]. sigmoidal [YC25]. signal [JJM25]. simplex [ZC25]. simulated [DG25]. singular [MS25]. SIS [KS25a]. SMEs [HBG25]. Solution [KIF^{+}25, KSC25]. solutions [BPT25, KC25]. solve [GLTW25]. solving [GHD25, HRST25, KC25, LGDN25, SQLZ25, SGRS25]. some [ZLS25]. sources [AAN25]. space [SHK25]. spaces [KIF^{+}25]. sparrow [HBG25]. sparse [JJM25, Sah25, WYZ25]. Spectral [TC25, GHD25, HRST25, ZMG25]. spline [LWL^{+}25]. splitting [CLJX25, WCS25]. stability [CCLH25, SX25]. stabilization [CCLH25]. stable [LLZH25]. standby [MHWY25]. state [GHZ25]. stepping [WX25]. Stochastic [MHWY25, CCLH25, KS25a]. strongly [KS25a, SHK25]. structural [ZLY25]. structured [Sah25]. study [UX25]. subdiffusion [YXM25]. Sufficient [Xie25]. supply [HBG25]. sure [CCLH25]. surrogate [WYZ25]. SVM [AGGM25]. switching [ZZ25]. Sylvester [BKST25]. symmetric [ZNL25]. system [DJ25, ZMG25]. systems [CCLH25, MHWY25, ZLY25].

Taylor [CA25]. Taylor-like [CA25]. technique [SWWY25]. tensor [BKST25, CWX25]. tensors [HLS25]. term [KSC25]. theorems [HLS25]. thermal [Sai25]. thermoelastic [Sai25]. thermoporoelasticity [ABH^{+}25]. third [GHD25, UX25]. third-grade [UX25]. three [LC25]. till [CM25, DJ25]. time [ABH^{+}25, CCLH25, KS25b, MHWY25, WX25, ZMG25]. time-stepping [WX25]. times [CP25]. Timestep [OTD^{+}25]. Timestep-dependent [OTD^{+}25]. transaction [AAN25]. transfer [ADY25]. transmission [TC25]. transport [JLC25]. triangles [ZBK25]. trust [Xie25]. two [CLJX25]. two-block [CLJX25]. type [BRPP25, CLJX25, SWWY25].

unbounded [ZMG25]. uncertain [LLG25]. unconditionally [LLZH25]. undamped [ZLY25]. uniform [MT25]. uniformly [DG25]. unreliable [MHWY25]. updating [ZLY25]. Using [BPT25, CA25, SSM25].

value [Ran25]. values [JFP25]. variable [MS25, Sai25]. via [BRPP25, KSC25, KIF^{+}25, ZC25]. viscoelastic [LWL^{+}25]. Volterra [AFA25, GHD25, MS25]. volume [PY25].

wave [LWL^{+}25]. wavelets [SKM25]. weakly [KIF^{+}25, MS25]. weighted [BRPP25]. white [Ran25]. without [JJM25].

zeros [FA25].

References

Alsenafi:2025:GAP

[AAN25] Abdulaziz Alsenafi, Fares Alazemi, and Alireza Najafi. Geometric Asian power option pricing with transaction cost under the geometric fractional Brownian motion with w sources of risk in fuzzy environment. Journal of Computational and Applied Mathematics, 453(??): ??, January 1, 2025. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (elec-

                                                   tronic).  URL http://www.
                                                   sciencedirect.com/science/
                                                   article/pii/S037704272400414X.

Alikhanov:2025:MMR

[ABH^{+}25] Anatoly Alikhanov, Huiran Bai, Jian Huang, Aleksei Tyrylgin, and Yin Yang. Multiscale model reduction for the time fractional thermoporoelasticity problem in fractured and heterogeneous media. Journal of Computational and Applied Mathematics, 455 (??):??, February 2025. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (elec-

                                                   tronic).  URL http://www.
                                                   sciencedirect.com/science/
                                                   article/pii/S0377042724004060.

Arias:2025:HQH

[ACO25] Luis A. Souto Arias, Pasquale Cirillo, and Cornelis W. Oosterlee. The Heston–Queue–Hawkes process: a new self-exciting jumpdiffusion model for options pricing, and an extension of the COS method for discrete distributions. Journal of Computational and Applied Mathematics, 454(??):??, January 15, 2025. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/S0377042724004266.

Ahmad:2025:FPA

[ADY25] Haroon Ahmad, Fahim Ud Din, and Mudasir Younis. A fixed point analysis of fractional dynamics of heat transfer in chaotic fluid layers. Journal of Computational and Applied Mathematics, 453(??):??, January 1, 2025. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/S0377042724003935.

Amirali:2025:SON

[AFA25] Ilhame Amirali, Burcu Fedakar, and Gabil M. Amiraliyev. Secondorder numerical method for a neutral Volterra integrodifferential equation. Journal of Computational and Applied Mathematics, 453(??): ??, January 1, 2025. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/S0377042724004096.

Astorino:2025:DCP

[AGGM25] Annabella Astorino, Manlio Gaudioso, Enrico Gorgone, and Benedetto Manca. Difference of convex programming in adversarial SVM. Journal of Computational and Applied Mathematics, 454(??):??, January 15, 2025. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (elec-

                                        tronic).  URL http://www.
                                        sciencedirect.com/science/
                                        article/pii/S0377042724004503.

Anonymous:2025:Ja

[Ano25a] Anonymous. 1 January 2025. Journal of Computational and Applied Mathematics, 453(??): ??, January 1, 2025. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic).

Anonymous:2025:Ma

[Ano25b] Anonymous. 1 March 2025. Journal of Computational and Applied Mathematics, 456(??):??, March 1, 2025. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic).

Anonymous:2025:Jb

[Ano25c] Anonymous. 15 January 2025. Journal of Computational and Applied Mathematics, 454(??): ??, January 15, 2025. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic).

Anonymous:2025:EBa

[Ano25d] Anonymous. Editorial Board. Journal of Computational and Applied Mathematics, 453(??): ??, January 1, 2025. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (elec-

                                        tronic).  URL http://www.
                                        sciencedirect.com/science/
                                        article/pii/S037704272400462X.

Anonymous:2025:EBb

[Ano25e] Anonymous. Editorial Board.

Journal of Computational and Applied Mathematics, 454(??): ??, January 15, 2025. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/S0377042724004801.

Anonymous:2025:EBc

[Ano25f] Anonymous. Editorial Board.

Journal of Computational and Applied Mathematics, 455(??): ??, February 2025. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/S037704272400493X.

Anonymous:2025:EBd

[Ano25g] Anonymous. Editorial Board.

Journal of Computational and Applied Mathematics, 456(??): ??, March 1, 2025. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/S0377042724005156.

Anonymous:2025:F

[Ano25h] Anonymous. February 2025.

Journal of Computational and Applied Mathematics, 455(??): ??, February 2025. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic).

Baz:2025:GMR

[BAPPF25] Juan Baz, Pedro Alonso, Juan Manuel Peña, and Raúl Pérez-Fernández. Gaussian Markov Random Fields over graphs of paths and high relative accuracy. Journal of Computational and Applied Mathematics, 453(??):??, January 1, 2025. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://www.

sciencedirect.com/science/
article/pii/S0377042724003911.

Basak:2025:EAS

[BC25] Arpita Basak and Amit Choudhury. Estimating arrival and service rates from queue length observations. Journal of Computational and Applied Mathematics, 455(??):??, February 2025. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/S037704272400445X.

Bagheri:2025:MBH

[BKST25] Mahsa Bagheri, Faranges Kyanfar, Abbas Salemi, and Azita Tajaddini. A modified block Hessenberg method for low-rank tensor Sylvester equation. Journal of Computational and Applied Mathematics, 456(??):??, March 1, 2025. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://www.

sciencedirect.com/science/
article/pii/S0377042724004588.

Boroujeni:2025:NSK

[BPT25] Ahmad Aliyari Boroujeni, Reza Pourgholi, and Seyed Hashem Tabasi. Numerical solutions of KDV and mKDV equations: Using sequence and multi-core parallelization implementation. Journal of Computational and Applied Mathematics, 454(??): ??, January 15, 2025. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/S0377042724004333.

Butt:2025:NCJ

[BRPP25] S. I. Butt, T. Rasheed, D. Pecarić, and J. Pecarić. New cyclic Jensen type bounds via weighted HH inequalities and Montgomery identity with application to information inequalities. Journal of Computational and Applied Mathematics, 453(??):??, January 1, 2025. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://www.

sciencedirect.com/science/
article/pii/S0377042724003868.

Chaskalovic:2025:EIA

[CA25] Joël Chaskalovic and Franck Assous. Enhancing interpolation and approximation error estimates using a novel Taylor-like formula. Journal of Computational and Applied Mathematics, 453(??):??, January 1, 2025. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (elec-

tronic).  URL http://www.

                                        sciencedirect.com/science/
                                        article/pii/S0377042724004023.

Cai:2025:ASE

[CCLH25] Ting Cai, Pei Cheng, Xing Liu, and Mingang Hua. Almost sure exponential stability and stochastic stabilization of discrete-time stochastic systems with impulses. Journal of Computational and Applied Mathematics, 453(??): ??, January 1, 2025. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (elec-

                                        tronic).  URL http://www.
                                        sciencedirect.com/science/
                                        article/pii/S0377042724004011.

Cuesta:2025:NAR

[CdlHG25] Carlota M. Cuesta, Francisco de la Hoz, and Ivan Girona. Numerical approximation of Riesz–Feller operators on R. Journal of Computational and Applied Mathematics, 454(??): ??, January 15, 2025. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (elec-

                                        tronic).  URL http://www.
                                        sciencedirect.com/science/
                                        article/pii/S0377042724004436.

Chao:2025:ERP

[CLJX25] Miantao Chao, Yongzi Lu, Jinbao Jian, and Xiao Xu. An efficient regularized PR splitting type algorithm for two-block nonconvex linear constrained programs in l_{1/2} regularized compressed sensing problems. Journal of Computational and Applied Mathematics, 453(??):??, January 1, 2025. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/S0377042724003947.

Cano:2025:EER

[CM25] B. Cano and M. J. Moreta.

Efficient exponential Rosenbrock methods till order four. Journal of Computational and Applied Mathematics, 453(??): ??, January 1, 2025. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/S0377042724004072.

Chadjiconstantinidis:2025:BDM

[CP25] Stathis Chadjiconstantinidis and Konstadinos Politis. Bounds for the distribution and moments of the forward and backward recurrence times in a renewal process. Journal of Computational and Applied Mathematics, 454 (??):??, January 15, 2025. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/S0377042724004151.

Cui:2025:GEM

[CQ25] Chunfeng Cui and Liqun Qi. A genuine extension of the Moore– Penrose inverse to dual matrices. Journal of Computational and Applied Mathematics, 454 (??):??, January 15, 2025. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (elec-

tronic).  URL http://www.

                                        sciencedirect.com/science/
                                        article/pii/S0377042724004345.

Cao:2025:DIH

[CWX25] Zhengbang Cao, Yimin Wei, and Pengpeng Xie. An L-DEIM induced high order tensor interpolatory decomposition. Journal of Computational and Applied Mathematics, 453(??): ??, January 1, 2025. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (elec-

                                        tronic).  URL http://www.
                                        sciencedirect.com/science/
                                        article/pii/S0377042724003923.

Diana:2025:NUS

[DG25] Derick Diana and Tim Gebbie. Non-uniformly sampled simulated price impact of an orderbook. Journal of Computational and Applied Mathematics, 456 (??):??, March 1, 2025. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (elec-

                                        tronic).  URL http://www.
                                        sciencedirect.com/science/
                                        article/pii/S0377042724004515.

Dembinska:2025:MLI

[DJ25] Anna Dembińska and Krzysztof Jasiński. Maximum likelihood inference about parameters of geometric lifetimes of heterogeneous components from data collected till failure of a k-out-of-n: G system. Journal of Computational and Applied Mathematics, 454(??):??, January 15, 2025. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (elec-

                                        tronic).  URL http://www.

sciencedirect.com/science/
article/pii/S0377042724004448.

Diegel:2025:ACD

[DLR25] Amanda E. Diegel, Xuejian Li, and Leo G. Rebholz. Analysis of continuous data assimilation with large (or even infinite) nudging parameters. Journal of Computational and Applied Mathematics, 456(??):??, March 1, 2025. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/S0377042724004709.

Frantsuzov:2025:GAB

[FA25] Viktor A. Frantsuzov and Anton V. Artemyev. A global argument-based algorithm for finding complex zeros and poles to investigate plasma kinetic instabilities. Journal of Computational and Applied Mathematics, 456(??):??, March 1, 2025. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/S0377042724004667.

Farouki:2025:AMC

[FKVZ25] Rida T. Farouki, Marjeta Knez, Vito Vitrih, and Emil Zagar. Application of a metric for complex polynomials to bounded modification of planar Pythagoreanhodograph curves. Journal of Computational and Applied Mathematics, 456(??):??, March 1, 2025. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://www.

                                        sciencedirect.com/science/
                                        article/pii/S0377042724004849.

Gidisu:2025:BDE

[GH25] Perfect Y. Gidisu and Michiel E. Hochstenbach. Block discrete empirical interpolation methods. Journal of Computational and Applied Mathematics, 454(??): ??, January 15, 2025. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (elec-

                                        tronic).  URL http://www.
                                        sciencedirect.com/science/
                                        article/pii/S0377042724004357.

Ghalini:2025:ESC

[GHD25] Rohollah Ghaedi Ghalini, Esmail Hesameddini, and Hojatollah Laeli Dastjerdi. An efficient spectral collocation method for solving Volterra delay integral equations of the third kind. Journal of Computational and Applied Mathematics, 454(??): ??, January 15, 2025. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (elec-

                                        tronic).  URL http://www.
                                        sciencedirect.com/science/
                                        article/pii/S037704272400387X.

Ge:2025:NEI

[GHZ25] Liang Ge, Jian Huang, and Jianwei Zhou. Novel efficient iterative schemes for linear finite element approximations of elliptic optimal control problems with integral constraint on the state. Journal of Computational and Applied Mathematics, 456(??):??, March 1, 2025. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://www.

sciencedirect.com/science/
article/pii/S0377042724004746.

Gong:2025:GBO

[GLTW25] Zhaohua Gong, Chongyang Liu, Kok Lay Teo, and Yonghong Wu. A gradient-based optimization algorithm to solve optimal control problems with conformable fractional-order derivatives. Journal of Computational and Applied Mathematics, 454 (??):??, January 15, 2025. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/S0377042724004187.

Hou:2025:ISS

[HBG25] Liangliang Hou, Gongbing Bi, and Qianqian Guo. An improved sparrow search algorithm optimized LightGBM approach for credit risk prediction of SMEs in supply chain finance. Journal of Computational and Applied Mathematics, 454(??): ??, January 15, 2025. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/S0377042724004461.

Ham:2025:FOF

[HKH^{+}25] Seokjun Ham, Seungyoon Kang, Youngjin Hwang, Gyeonggyu Lee, Soobin Kwak, Jyoti, and Junseok Kim. A fourthorder finite difference method for the Allen–Cahn equation. Journal of Computational and Applied Mathematics, 453(??): ??, January 1, 2025. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (elec-

                                        tronic).  URL http://www.
                                        sciencedirect.com/science/
                                        article/pii/S0377042724004084.

He:2025:PEI

[HLS25] Jun He, Yanmin Liu, and Xiaowei Shen. Pareto eigenvalue inclusion theorems and copositive tensors. Journal of Computational and Applied Mathematics, 454(??):??, January 15, 2025. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (elec-

                                        tronic).  URL http://www.
                                        sciencedirect.com/science/
                                        article/pii/S0377042724004424.

Huang:2025:SCS

[HRST25] Yu Huang, Narges Tohidi Rad, Mohammad Hadi Noori Skandari, and Emran Tohidi. A spectral collocation scheme for solving nonlinear delay distributedorder fractional equations. Journal of Computational and Applied Mathematics, 456(??):??, March 1, 2025. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://www.

                                        sciencedirect.com/science/
                                        article/pii/S037704272400476X.

Jafari:2025:ALP

[JFP25] A. A. Jafari, S. Firoozi, and A. Pak. Analysis of the lifetime performance index in Gompertz distribution based on krecord values. Journal of Computational and Applied Mathematics, 453(??):??, January 1, 2025. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/S037704272400400X.

Jian:2025:IBA

[JJM25] Jinbao Jian, Jiachen Jin, and Guodong Ma. An IDFPM-based algorithm without Lipschitz continuity to constrained nonlinear equations for sparse signal and blurred image restoration problems. Journal of Computational and Applied Mathematics, 456 (??):??, March 1, 2025. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/S0377042724004837.

Jiang:2025:DGM

[JLC25] Xia Jiang, Rui Li, and Zhangxin Chen. A discontinuous Galerkin method for the Brinkman– Darcy-transport problem. Journal of Computational and Applied Mathematics, 453(??): ??, January 1, 2025. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/S0377042724004047.

Jacob:2025:CQN

[JVW25] Birgit Jacob, Lukas Vorberg, and Christian Wyss. Computing the quadratic numerical range. Journal of Computational and Applied Mathematics, 453 (??):??, January 1, 2025. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (elec-

                                        tronic).  URL http://www.
                                        sciencedirect.com/science/
                                        article/pii/S0377042724002991.

Karageorghis:2025:MLM

[KC25] Andreas Karageorghis and C. S. Chen. Multi-level method of fundamental solutions for solving polyharmonic problems. Journal of Computational and Applied Mathematics, 456(??):??, March 1, 2025. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://www.

                                        sciencedirect.com/science/
                                        article/pii/S0377042724004692.

Kumar:2025:AFR

[KCM25] D. Kumar, A. K. B. Chand, and P. R. Massopust. Approximation with fractal radial basis functions. Journal of Computational and Applied Mathematics, 454(??):??, January 15, 2025. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (elec-

                                        tronic).  URL http://www.
                                        sciencedirect.com/science/
                                        article/pii/S0377042724004497.

Kattan:2025:SIE

[KIF^{+}25] Doha A. Kattan, Umar Ishtiaq, Muhammad Farhan, Ioannis K. Argyros, Mohammad Akram, and Samundra Regmi. Solution of an integral equation in controlled rectangular metric spaces via weakly contractive mappings.

Journal of Computational and Applied Mathematics, 454(??): ??, January 15, 2025. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/S0377042724004369.

Kiouvrekis:2025:DPS

[KS25a] Yiannis Kiouvrekis and Ioannis S.

Stamatiou. Domain preserving and strongly converging explicit scheme for the stochastic SIS epidemic model. Journal of Computational and Applied Mathematics, 456(??):??, March 1, 2025. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/S0377042724004680.

Koellermeier:2025:PIM

[KS25b] Julian Koellermeier and Giovanni Samaey. Projective integration methods in the Runge– Kutta framework and the extension to adaptivity in time. Journal of Computational and Applied Mathematics, 454(??): ??, January 15, 2025. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/S0377042724003960.

Katani:2025:ASM

[KSC25] Roghayeh Katani, Sedaghat Shahmorad, and Dajana Conte. Approximate solution of multiterm fractional differential equations via a block-by-block method. Journal of Computational and Applied Mathematics, 453(??): ??, January 1, 2025. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/S0377042724003844.

Lovat:2025:LDE

[LC25] Giampiero Lovat and Salvatore Celozzi. Lucas decomposition and extrapolation methods for the evaluation of infinite integrals involving the product of three Bessel functions of arbitrary order. Journal of Computational and Applied Mathematics, 453(??):??, January 1, 2025. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/S037704272400390X.

Linghu:2025:HOM

[LGDN25] Jiale Linghu, Weifeng Gao, Hao Dong, and Yufeng Nie. Higher-order multi-scale physicsinformed neural network (HOMS-PINN) method and its convergence analysis for solving elastic problems of authentic composite materials. Journal of Computational and Applied Mathematics, 456(??):??, March 1, 2025. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/S0377042724004722.

Liu:2025:MLE

[LLG25] Zhe Liu, Yang Liu, and Rong Gao. Maximum likelihood estimation for multi-factor uncertain differential equations. Journal of Computational and Applied Mathematics, 454(??): ??, January 15, 2025. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/S0377042724004540.

Li:2025:SOL

[LLZH25] Yaoda Li, Min Li, Boya Zhou, and Zhibin Han. A second-order linear unconditionally energy-stable scheme for the phase field crystal equation. Journal of Computational and Applied Mathematics, 455 (??):??, February 2025. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/S0377042724004527.

Li:2025:CBS

[LWL^{+}25] Yaomeng Li, Feng Wang, Qiao Li, Chao Fu, and Xu Guo. Cubic B-spline based elastic and viscoelastic wave propagation method. Journal of Computational and Applied Mathematics, 456(??):??, March 1, 2025. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/S0377042724004850.

Ma:2025:SMO

[MHWY25] Mengrao Ma, Linmin Hu, Lin Wang, and Xiaoyun Yu. Stochastic modeling and optimization of discrete-time cold standby repairable systems with unreliable repair facility and retrial mechanism. Journal of Computational and Applied Mathematics, 453(??):??, January 1, 2025. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (elec-

                                        tronic).  URL http://www.
                                        sciencedirect.com/science/
                                        article/pii/S0377042724003832.

Masirevic:2025:ISR

[MPB25] Dragana Jankov Masirević, Tibor K. Pogány, and Tomislav Burić. Integral and series representations of the Nuttall function. Journal of Computational and Applied Mathematics, 453 (??):??, January 1, 2025. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (elec-

                                        tronic).  URL http://www.
                                        sciencedirect.com/science/
                                        article/pii/S0377042724003881.

Ma:2025:NMB

[MS25] Zheng Ma and Martin Stynes. A Nyström method based on product integration for weakly singular Volterra integral equations with variable exponent. Journal of Computational and Applied Mathematics, 454(??): ??, January 15, 2025. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (elec-

                                        tronic).  URL http://www.

sciencedirect.com/science/
article/pii/S0377042724004138.

Masoodi:2025:OUS

[MT25] Mehdi Masoodi and Haniyeh Taromideh. Optimal and uniform sensor arrangement in nearfield imaging. Journal of Computational and Applied Mathematics, 454(??):??, January 15, 2025. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/S0377042724004370.

Novo:2025:CNN

[NT25] Julia Novo and Eduardo Terrés.

Can neural networks learn finite elements? Journal of Computational and Applied Mathematics, 453(??):??, January 1, 2025. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/S0377042724004175.

Orynyak:2025:TDE

[OTD^{+}25] Igor Orynyak, Anton Tsybulnyk, Kirill Danylenko, Andrii Oryniak, and Sergii Radchenko. Timestep-dependent element interpolation functions in the method of matched sections on the example of heat conduction problem. Journal of Computational and Applied Mathematics, 456(??):??, March 1, 2025. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (elec-

tronic).  URL http://www.

sciencedirect.com/science/
article/pii/S0377042724004710.

Ozarslan:2025:AFC

[Öza25] Mehmet Ali Özarslan. On the approximation to fractional calculus operators with multivariate Mittag-Leffler function in the kernel. Journal of Computational and Applied Mathematics, 454(??):??, January 15, 2025. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/S0377042724003972.

Peng:2025:ALE

[PY25] Gang Peng and Di Yang. An arbitrary Lagrangian–Eulerian positivity-preserving finite volume scheme for radiation hydrodynamics equations in the equilibrium-diffusion limit. Journal of Computational and Applied Mathematics, 455(??): ??, February 2025. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/S0377042724004059.

Ranjbar:2025:EES

[Ran25] Hassan Ranjbar. An explicit Euler scheme for generalized n-dimensional second-order differential equations with initial value conditions driven by additive Gaussian white noises. Journal of Computational and Applied Mathematics, 454(??): ??, January 15, 2025. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/S037704272400428X.

Saha:2025:IEP

[Sah25]  Tanay Saha.  Inverse eigen-

value problem for affinely structured or sparse polynomial matrices with prescribed eigenvalues. Journal of Computational and Applied Mathematics, 454 (??):??, January 15, 2025. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/S0377042724004564.

Said:2025:IGR

[Sai25] Samia M. Said. Influence of gravitational on a rotating nonlocal thermoelastic medium with thermal variable conductivity. Journal of Computational and Applied Mathematics, 454(??): ??, January 15, 2025. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/S0377042724004291.

Singh:2025:HEO

[SGRS25] Randhir Singh, Vandana Guleria, Higinio Ramos, and Mehakpreet Singh. Highly efficient optimal decomposition approach and its mathematical analysis for solving fourth-order Lane– Emden–Fowler equations. Journal of Computational and Applied Mathematics, 456(??):??, March 1, 2025. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://www.

                                        sciencedirect.com/science/
                                        article/pii/S0377042724004874.

Shadimetov:2025:OQF

[SHK25] Kholmat Shadimetov, Abdullo Hayotov, and Umedjon Khayriev. Optimal quadrature formulas for approximating strongly oscillating integrals in the Hilbert space W̃_{2}(m, m − 1) of periodic functions. Journal of Computational and Applied Mathematics, 453 (??):??, January 1, 2025. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (elec-

                                        tronic).  URL http://www.
                                        sciencedirect.com/science/
                                        article/pii/S0377042724003820.

Singh:2025:EWM

[SKM25] Akanksha Singh, Ankur Kanaujiya, and Jugal Mohapatra. Euler wavelets method for optimal control problems of fractional integro-differential equations. Journal of Computational and Applied Mathematics, 454 (??):??, January 15, 2025. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (elec-

                                        tronic).  URL http://www.
                                        sciencedirect.com/science/
                                        article/pii/S0377042724004278.

Shen:2025:LGM

[SNZ25] Binrui Shen, Qiang Niu, and Shengxin Zhu. Lightning graph matching. Journal of Computational and Applied Mathematics, 454(??):??, January 15, 2025. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/S0377042724004382.

Shang:2025:NPF

[SQLZ25] Youlin Shang, Deqiang Qu, Junxiang Li, and Roxin Zhang. A new parameter-free continuously differentiable filled function algorithm for solving nonlinear equations and data fitting problems. Journal of Computational and Applied Mathematics, 454 (??):??, January 15, 2025. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/S0377042724004473.

Spiridonov:2025:PDO

[SSM25] Denis Spiridonov, Sergei Stepanov, and Tina Mai. Prediction of discretization of online GMsFEM using deep learning for Richards equation. Journal of Computational and Applied Mathematics, 454(??):??, January 15, 2025. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/S0377042724004163.

Song:2025:EAT

[SWWY25] Haiming Song, Hao Wang, Jiageng Wu, and Jinda Yang. An efficient ADAM-type algorithm with finite elements discretization technique for random elliptic optimal control problems. Journal of Computational and Applied Mathematics, 454(??): ??, January 15, 2025. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (elec-

                                        tronic).  URL http://www.
                                        sciencedirect.com/science/
                                        article/pii/S0377042724004485.

Shah:2025:ASC

[SX25] Syed Omar Shah and Yonghui Xia. An analysis on the stability and controllability of nonlinear measure differential equations. Journal of Computational and Applied Mathematics, 453 (??):??, January 1, 2025. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (elec-

                                        tronic).  URL http://www.
                                        sciencedirect.com/science/
                                        article/pii/S0377042724003893.

Tan:2025:SAE

[TC25] Ting Tan and Waixiang Cao. Spectral approximation and error analysis for the transmission eigenvalue problem with an isotropic inhomogeneous medium. Journal of Computational and Applied Mathematics, 453(??):??, January 1, 2025. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (elec-

                                        tronic).  URL http://www.
                                        sciencedirect.com/science/
                                        article/pii/S0377042724004126.

Ullah:2025:CSH

[UX25] Inayat Ullah and Yingxiang Xu. Computational study of homogeneous–heterogeneous reactions in bioconvective thirdgrade material flow. Journal of Computational and Applied Mathematics, 454(??): ??, January 15, 2025. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/S0377042724004308.

Wang:2025:IMB

[WCS25] Lu-Xin Wang, Yang Cao, and Qin-Qin Shen. Improved modulus-based matrix splitting iteration methods for a class of horizontal implicit complementarity problems. Journal of Computational and Applied Mathematics, 454(??):??, January 15, 2025. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://www.

sciencedirect.com/science/
article/pii/S0377042724004321.

Wang:2025:CEH

[WX25] Weiwen Wang and Chuanju Xu. A class of efficient highorder time-stepping methods for the anisotropic phase-field dendritic crystal growth model. Journal of Computational and Applied Mathematics, 453(??): ??, January 1, 2025. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/S0377042724004102.

Wang:2025:SKM

[WYZ25] Ze Wang, Jun-Feng Yin, and Ji-Chen Zhao. The sparse Kaczmarz method with surrogate hyperplane for the regularized basis pursuit problem. Journal of Computational and Applied Mathematics, 454(??): ??, January 15, 2025. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/S037704272400431X.

Xie:2025:SCE

[Xie25] Pengcheng Xie. Sufficient conditions for error distance reduction in the l^{2} -norm trust region between minimizers of local nonconvex multivariate quadratic approximates. Journal of Computational and Applied Mathematics, 453(??):??, January 1, 2025. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/S0377042724003959.

Yu:2025:CAR

[YC25] Dansheng Yu and Feilong Cao. Construction and approximation rate for feedforward neural network operators with sigmoidal functions. Journal of Computational and Applied Mathematics, 453(??):??, January 1, 2025. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/S0377042724003996.

Yousef:2025:PBL

[You25] Soleiman Yousef. A posterioribased, local multilevel mesh refinement for the Darcy porous media flow problem. Journal of Computational and Applied Mathematics, 454(??): ??, January 15, 2025. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/S0377042724004114.

Yao:2025:EAM

[YWW25] Guoqing Yao, Zicheng Wang, and Zhongqing Wang. An efficient and accurate mapping method for elliptic equations in irregular annular domains. Journal of Computational and Applied Mathematics, 456(??):??, March 1, 2025. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://www.

sciencedirect.com/science/
article/pii/S0377042724004862.

Yan:2025:BIN

[YXM25] Xiong-Bin Yan, Zhi-Qin John Xu, and Zheng Ma. Bayesian Inversion with Neural Operator (BINO) for modeling subdiffusion: Forward and inverse problems. Journal of Computational and Applied Mathematics, 454(??):??, January 15, 2025. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/S0377042724004400.

Zhou:2025:NQG

[ZBK25] Jun Zhou, Pieter J. Barendrecht, and Jirí Kosinka. Numerical quadrature for Gregory triangles. Journal of Computational and Applied Mathematics, 453 (??):??, January 1, 2025. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/S0377042724003984.

Zhou:2025:QHO

[ZC25] Yongxiong Zhou and Ruyun Chen. On quadrature of highly oscillatory Bessel function via asymptotic analysis of simplex integrals. Journal of Computational and Applied Mathematics, 456(??):??, March 1, 2025. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/S0377042724004886.

Zhukova:2025:ANA

[ZFCM25] Aleksandra Zhukova, Anna Flerova, Alexey Chernov, and Georgy Milyuchikhin. Approaches to numerical analysis of optimal control with linear phase constraints on the example of the assets and liabilities management by a bank. Journal of Computational and Applied Mathematics, 453(??):??, January 1, 2025. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/S0377042724003790.

Zhang:2025:DIE

[ZLS25] Zecheng Zhang, Wing Tat Leung, and Hayden Schaeffer. A discretization-invariant extension and analysis of some deep operator networks. Journal of Computational and Applied Mathematics, 456(??):??, March 1, 2025. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/S0377042724004758.

Zeng:2025:IMU

[ZLY25] Min Zeng, Xianlu Liao, and Yongxin Yuan. An iterative method for updating undamped structural systems with connectivity constraints. Journal of Computational and Applied Mathematics, 454(??): ??, January 15, 2025. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/S0377042724004394.

Zhang:2025:ECS

[ZMG25] Xiaohao Zhang, Liquan Mei, and Shimin Guo. Energy-conserving SAV–Hermite–Galerkin spectral scheme with time adaptive method for coupled nonlinear Klein–Gordon system in multidimensional unbounded domains. Journal of Computational and Applied Mathematics, 454(??): ??, January 15, 2025. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (elec-

tronic).  URL http://www.

                                        sciencedirect.com/science/
                                        article/pii/S0377042724004539.

Zhu:2025:EMS

[ZNL25] Hong Zhu, Chenchen Niu, and Yongjin Liang. Efficient method for symmetric nonnegative matrix factorization with an approximate augmented Lagrangian scheme. Journal of Computational and Applied Mathematics, 454(??):??, January 15, 2025. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (elec-

                                        tronic).  URL http://www.
                                        sciencedirect.com/science/
                                        article/pii/S0377042724004679.

Zhou:2025:MIF

[ZWSC25] Sheng-Wei Zhou, Chun Wen, Zhao-Li Shen, and Bruno Carpentieri. The MFPIO iteration and the FPMPE method for multilinear PageRank computations. Journal of Computational and Applied Mathematics, 454(??):??, January 15, 2025. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (elec-

                                        tronic).  URL http://www.
                                        sciencedirect.com/science/
                                        article/pii/S0377042724004412.

Zheng:2025:GIE

[ZZ25] Yawen Zheng and Song-Ping Zhu. A generalized integral equation formulation for pricing American options under regime-switching model. Journal of Computational and Applied Mathematics, 453(??):??, January 1, 2025. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/S0377042724002668.